Use elimination to solve the system of equations.
4x-9y=15
3x-4y=14

Answer by checkley77(12844)

4x-9y=15 multiply this equation by -3
3x-4y=14 multiply thisequation by 4 & then add them.
-12x+27y=-45
12x-16y=56
------------------
11y=11
y=11/11
y=1 answer.
4x-9*1=15
4x-9=15
4x=15+9
4x=24
x=24/4
x=6 answer.
Proof
3*6-4*1=14
18-4=14
14=14